### v4

# Projected Augmented Waves (PAW): extended resolution of unity method

Abstract
The Projected Augmented Waves (PAW) method is based on a linear transformation between the pseudo wavefunctions and the all electron wavefunctions. To obtain high accuracy with this method, it is important that the local part of the linear transform (inside each atomic sphere) be defined over a complete basis set (with deviations from completeness leading to corrections to the total energy not computed within current implementations of PAW). Here we show how to make this basis much closer to complete without significant additional computational work and without modifying the transformation in any significant way thereby making the modifications we propose easy to implement in current electronic structure codes for PAW.. This is done by extending the resolution of unity used for the transform to include more smooth wavefunctions (which have nothing to do with the atomic problem) and having them linearly transform via the identity.
I Introduction
It is well know that all electron methods are often more technically difficult to practically implement then pseudopotential methods [3, 8, 7, 13, 12, 1, 11, 9, 10, 2, 14, 6, 5, 4]. This is in part due to the complex nature of the basis wavefunctions, for all electron methods, which are often based on augmentation methods - where the wavefunction has markedly different expressions in the Muffin Tin (MT) regions and in the interstitial regions respectively. This augmentation construction often complicates the evaluation of the Khon Sham (KS) Hamiltonian and related quantities needed to solve electronic structure problems. This leads to some technical disadvantages of all electron methods vs. pseudopotential ones, in particular:
  • For all electron methods the calculation of eigenvalues and eigenvectors of the Khon Sham (KS) Hamiltonian often requires FLAPW (Full Potential Linearized Augmented Plane Waves [22, 16, 18, 21, 15, 19, 20, 17]) which requires a relatively complex setup with multiple steps and calculations. On the other hand full potential plane wave methods have simple expressions for KS Hamiltonian matrix elements and reduce to the identity matrix for wavefunction overlaps in many cases - in particular for norm conserving pseudopotentials.
  • The Coulomb term (Hartree energy) is more complex in augmented, all electron, systems. Indeed it requires the Weinert method for efficient calculation and convergence without shape approximations [23, 24]. Where as for pseudopotentials Ewald sums suffice for most applications [25].
  • Atomic force calculations in all electron methods are more complex as they require the computation of the Pulay contribution to the forces - which arise due to changes of span of the all electron basis set with atomic positions (the Hellman-Feynman theorem does not apply directly in this case [26] and needs modifications), while many pseudized methods require only the Hellman-Feynman terms.
  • Substitutional alloys are difficult to study for all electron methods (other then through cumbersome methods such as Korringa Khon Rostoker (KKR) methods [28, 27] or through supercells), where as for pseudopotentials there are many viable virtual crystal approximation methods [29, 30].
  • The Car Parrinello molecular dynamics simulations [31, 25, 32, 33] method is significantly easier with pseudopotentials then with augmented waves such as Linearized Augmented Plane Waves (LAPW) and very limited work has been done so far with all electron methods in that direction [25, 1].
As such, despite many advances in the accuracy and efficiency of all electron methods and their conceptual simplicity [36, 35, 34, 11, 12, 10, 9, 7, 8, 37], it is extremely worthwhile to study pseudopotential methods (atleast currently) - as they are currently the leading methods for electronic structure calculations within Density Functional Theory (DFT). Most modern pseudopotential methods may be loosely divided into three different categories:
  • Norm conserving pseudopotentials [2, 25]
  • Ultrasoft psudopotentials [38, 2]
  • Projected Augmented Waves (PAW) [39, 40, 41]
We now describe each of these three electronic structure methods in turn.
Norm conserving pseudopotentials are based on the idea that a perfectly spherical scattering potential (the atomic sphere, when small enough, is roughly spherically symmetric) can be represented by a scattering phase shift - in each partial wave and at each energy - as a boundary condition at the surfaces of the sphere. Any potential producing these scattering phase shifts is sufficient for calculations of eigenvalues of the KS Hamiltonian - not just the original atomic one. Therefore an equivalent problem where a much smoother potential with the same scattering phase shift (in the middle of the valence band) as the atomic potential, in each different angular momentum channel, can be used in the solution of the KS Hamiltonian. These smoother potentials may be used to solve for the eigenvalues of the KS problem leading to lower cutoffs (smoother wavefunctions) needed for electronic structure calculations. The norm conserving condition insures both accurate Madelung energy and that the derivative of the scattering phase shift with respect to the energy is the same both for the atomic potential and the pseudopotential leading to improved accuracy as a function of energy (better transferability).
A competitor to norm conserving pseudopotentials are ultrasoft potentials. These pseudopotentials use the many state generalizations of the Kleinman-Bylander transform [2] in order to reproduce the scattering phase shifts of the atomic problem simultaneously at many energies. This allows one to get rid of the norm conserving condition (needed for accuracy and transferability without multiple scattering energies) and restore correct Madelung energies and with a fictitious overlap matrix between the pseudized wavefunctions and compensating charges (for higher multipoles). This greatly softens the potential as the pseudopotential need not conserve norm and therefore has fewer constraints leading to softer pseudopotentials, e.g. especially for $3d$ transition metals which are hard to soften while conserving norm as the $3d$ wavefunctions have no oscillations to smooth out with a pseudized wavefunction [25, 2, 38]. This also solves the problem of ghost bands: semi-core bands with significantly different scattering phase shifts then in the middle of the valence band that often plagued norm conserving pseudopotentials can now be handled by introducing a Kleinman-Bylander transform with a state at the energy of the semi-core states thereby insuring good wavefunctions there. The price for this is minor and associated with some additional terms in the KS equations which come about because of the fictitious overlap (norm).
For PAW we introduce a linear transformation between the soft (low cutoff) pseudized wavefunctions and the all electron wavefunctions. The various quantities needed to set up the KS equation and obtain the ground state energy are then expressed through this linear transformation. PAW is sometimes considered the most versatile because at its core its an efficient method to map a pseudo-wave function problem onto an all electron problem - leading to great accuracy of computations. One of its limitations is that it assumes some completeness relations for pseudo wavefunctions used in the transform, a problem we here partially remedy at very low computational cost and in a manner that is highly compatible with current electronic structure codes for PAW.
II PAW Summary
We now describe PAW in detail. First we note that the all electron wavefunction is given through the relationship:
$\psi=\mathcal{T}\tilde{\psi}$ (1)
where $\tilde{\psi}$ is the a pseudo wavefunction. We now write:
$\mathcal{T}=\mathbb{I}+\sum_{\mathbf{R}}S_{\mathbf{R}}$ (2)
Here $\mathbf{R}$ are the sites of the atomic nuclei and $S_{\mathbf{R}}$ is a linear transformation localized at the augmentation sphere centered at $\mathbf{R}$. More precisely:
$S_{\mathbf{R}}=\sum_{i\in\mathbf{R}}\left[\left|\phi_{i}\right\rangle-\left|\tilde{\phi}_{i}\right\rangle\right]\left\langle\tilde{p}_{i}\right|$ (3)
Where $\left|\phi_{i}\right\rangle$ are from atomic calculations representing the exact atomic states while $\left|\tilde{\phi}_{i}\right\rangle$ are smooth functions localized inside the augmentation spheres. Where:
$\left\langle\tilde{p}_{i}\mid\tilde{\phi}_{j}\right\rangle=\delta_{ij}$ (4)
There are many options for the wavefunctions $\left|\tilde{p}_{i}\right\rangle$, for example:
$\left|\tilde{p}_{i}\right\rangle=\sum_{j\in\mathbf{R}}\left|\tilde{\phi}_{j}\right\rangle\left[M\right]_{ji}^{-1}$ (5)
In this work we will consider only this case (though similar considerations may be made for more general cases [40]). Where
$M_{ij}=\left\langle\tilde{\phi}_{i}\mid\tilde{\phi}_{j}\right\rangle$ (6)
This means that
$\left[\sum_{i\in\mathbf{R}}\left|\tilde{\phi}_{i}\right\rangle\left\langle\tilde{p}_{i}\right|\right]\left|\tilde{\phi}_{j}\right\rangle=\left|\tilde{\phi}_{j}\right\rangle$ (7)
so that
$\sum_{i\in\mathbf{R}}\left|\tilde{\phi}_{i}\right\rangle\left\langle\tilde{p}_{i}\right|=\mathbb{I}_{\mathbf{R}}$ (8)
provided that the $\left|\tilde{\phi}_{i}\right\rangle$ form a complete basis set. However we have that in reality:
$\sum_{i\in\mathbf{R}}\left|\tilde{\phi}_{i}\right\rangle\left\langle\tilde{p}_{i}\right|=\left\{\begin{array}[]{cc}\mathbb{I}_{\mathbf{R}}&\in Span\left\{\left|\tilde{\phi}_{i}\right\rangle\right\}\\ 0&otherwise\end{array}\right.$ (9)
This relation, assumed completeness Eq. (8), is used to prove the identity that for a local operator:
$\left\langle\tilde{\psi}\right|\hat{A}\left|\tilde{\psi}\right\rangle_{\mathbf{R}}=\left\langle\tilde{\psi}_{\mathbf{R}}\right|\hat{A}\left|\tilde{\psi}_{\mathbf{R}}\right\rangle$ (10)
Where $\left\langle\right\rangle_{\mathbf{R}}$ means all integrations are done over the atomic sphere centered about $\mathbf{R}$. Where:
$\left|\psi_{\mathbf{R}}\right\rangle=\left[\sum_{i\in\mathbf{R}}\left|\phi_{i}\right\rangle\left\langle\tilde{p}_{i}\right|\right]\left|\tilde{\psi}\right\rangle$ (11)
These relations are used to compute the form of the density of electrons the correlation and exchange and the Hartree (Coulomb piece) of the Hamiltonian [39, 2, 25]. This is essential to current PAW methodology [40]. Here we improve on this resolution of identity without changing its mathematical form or incurring significant additional computational costs.
III Main Construction
We now address the main problem pointed out in the previous section that the $\left|\tilde{\phi}_{i}\right\rangle$ do not form a complete basis set. This is in particular because the set $\left|\tilde{\phi}_{i}\right\rangle$ of the same cardinality as $\left|\phi_{i}\right\rangle$ and we must obtain $\left|\phi_{i}\right\rangle$ from atomic calculations, where typically only two or three wavefunctions, per angular momentum channel, are available with current PAW atomic data sets [40]. Here we propose to circumvent this bottleneck by introducing:
$\psi=\mathcal{T}_{M}\tilde{\psi}$ (12)
where $\tilde{\psi}$ is the a pseudo wavefunction and $M$ stands for multiple states. We now write:
$\mathcal{T}_{M}=\mathbb{I}+\sum_{\mathbf{R}}S_{\mathbf{R}}^{M}$ (13)
Where
$S_{\mathbf{R}}^{M}=\sum_{i,\alpha\in\mathbf{R}}\left[\left|\phi_{i\alpha}\right\rangle-\left|\tilde{\phi}_{i\alpha}\right\rangle\right]\left\langle\tilde{p}_{i\alpha}\right|$ (14)
Where
$\left|\phi_{i\alpha}\right\rangle=\left\{\begin{array}[]{cc}\left|\phi_{i}\right\rangle&\alpha=1\\ \left|\tilde{\phi}_{i\alpha}\right\rangle&\alpha>1\end{array}\right.$ (15)
Here $\alpha$ runs over a list of on the order of ten to one hundred additional pseudo wavefunctions per an all electron wavefunction. Where we have used the idea that it is much easier to generate smooth (low cutoff) wave functions - say by a combination of Bessel functions with different energies multiplied by spherical harmonics with different angular momentum channels, then to do atomic data set calculations. These wavefunctions can be chosen to have good overlaps with the plane waves (Bessel functions) used for the basis for the KS problem for PAW in various angular momentum channels and therefore can be chosen as a nearly complete basis set for the space of the pseudized wave functions considered in the PAW method. As such multiple smooth wavefunctions can easily be added to the same all electron atomic calculation wavefunction. Where we have that:
$\left\langle\tilde{p}_{i\alpha}\mid\tilde{\phi}_{j\beta}\right\rangle=\delta_{ij}\delta_{\alpha\beta}$ (16)
There are many options for the wavefunctions $\left|\tilde{p}_{i\alpha}\right\rangle$ for example:
$\left|\tilde{p}_{i\alpha}\right\rangle=\sum_{j\in\mathbf{R}}\left|\tilde{\phi}_{j\beta}\right\rangle\left[M\right]_{j\beta i\alpha}^{-1}$ (17)
Where
$M_{i\alpha j\beta}=\left\langle\tilde{\phi}_{i\alpha}\mid\tilde{\phi}_{j\beta}\right\rangle$ (18)
We notice that:
$\sum_{i\in\mathbf{R}}\left|\tilde{\phi}_{i\alpha}\right\rangle\left\langle\tilde{p}_{i\alpha}\right|=\left\{\begin{array}[]{cc}\mathbb{I}_{\mathbf{R}}&\in Span\left\{\left|\tilde{\phi}_{i\alpha}\right\rangle\right\}\\ 0&otherwise\end{array}\right.$ (19)
So is much closer to the resolution of identity, as such we have inserted the identity transform in a clever way improving the correctness of the resolution of identity.
IV Some technical details
In this section we present the main technical consequences of our new transform as the apply to current implementations of PAW. However we note that since the mathematical structUre of our resolution of identity is identical to the one usually used for PAW these technical consequences would already be coded in electronic structure codes for PAW and require no further work to implement.
IV.1 Main modifications to local operators
We now write:
$\left|\psi_{\mathbf{R}}^{M}\right\rangle$ $=\left[\sum_{i\alpha\in\mathbf{R}}\left|\phi_{i\alpha}\right\rangle\left\langle\tilde{p}_{i\alpha}\right|\right]\left|\tilde{\psi}\right\rangle$
$\left|\tilde{\psi}_{\mathbf{R}}^{M}\right\rangle$ $=\left[\sum_{i\alpha\in\mathbf{R}}\left|\tilde{\phi}_{i\alpha}\right\rangle\left\langle\tilde{p}_{i\alpha}\right|\right]\left|\tilde{\psi}\right\rangle$ (20)
Now for a local operator $\hat{A}$ we write:
$\left\langle\mathcal{T}_{M}^{\dagger}\hat{A}\mathcal{T}_{M}\right\rangle$
$=\left[\left\langle\tilde{\psi}\right|+\sum_{\mathbf{R}}\left[\left\langle\psi_{\mathbf{R}}^{M}\right|-\left\langle\tilde{\psi}_{\mathbf{R}}^{M}\right|\right]\right]\times$
×$\times\hat{A}\left[\left|\tilde{\psi}\right\rangle+\sum_{\mathbf{R}}\left[\left|\psi_{\mathbf{R}}^{M}\right\rangle-\left|\tilde{\psi}_{\mathbf{R}}^{M}\right\rangle\right]\right]$
$=\left[\left\langle\tilde{\psi}\right|-\sum_{\mathbf{R}}\left\langle\tilde{\psi}_{\mathbf{R}}^{M}\right|\right]\hat{A}\left[\left|\tilde{\psi}\right\rangle-\sum_{\mathbf{R}^{\prime}}\left|\tilde{\psi}_{\mathbf{R}^{\prime}}^{M}\right\rangle\right]+$
$+\sum_{\mathbf{R}}\left\langle\psi_{\mathbf{R}}^{M}\right|\hat{A}\left|\psi_{\mathbf{R}}^{M}\right\rangle$
$=\left\langle\tilde{\psi}\right|\hat{A}\left|\tilde{\psi}\right\rangle+\sum_{\mathbf{R}}\left[\left\langle\psi_{\mathbf{R}}^{M}\right|\hat{A}\left|\psi_{\mathbf{R}}^{M}\right\rangle-\left\langle\tilde{\psi}_{\mathbf{R}}^{M}\right|\hat{A}\left|\tilde{\psi}_{\mathbf{R}}^{M}\right\rangle\right]$ (21)
It is now also useful to introduce:
$D_{i\alpha j\beta}^{\mathbf{R}}=\sum_{n,\mathbf{k}}f\left(\varepsilon_{n}\left(\mathbf{k}\right)\right)\left\langle\tilde{\psi}_{n}\left(\mathbf{k}\right)\mid\tilde{p}_{i\alpha}\right\rangle\left\langle\tilde{p}_{j\beta}\mid\tilde{\psi}_{n}\left(\mathbf{k}\right)\right\rangle$ (22)
This means that:
$\left\langle\mathcal{T}_{M}^{\dagger}\hat{A}\mathcal{T}_{M}\right\rangle$
$=\sum_{n,\mathbf{k}}f\left(\varepsilon_{n}\left(\mathbf{k}\right)\right)\left\langle\tilde{\psi}_{n}\left(\mathbf{k}\right)\right|\hat{A}\left|\tilde{\psi}_{n}\left(\mathbf{k}\right)\right\rangle+$
$+\sum_{\mathbf{R},i,j}D_{i\alpha j\beta}^{\mathbf{R}}\left[\left\langle\phi_{i\alpha}\right|\hat{A}\left|\phi_{j\beta}\right\rangle-\left\langle\tilde{\phi}_{i\alpha}\right|\hat{A}\left|\tilde{\phi}_{j\beta}\right\rangle\right]$
$+\sum_{core\in\mathbf{R}}\left\langle\phi_{c,i}\right|\hat{A}\left|\phi_{c,i}\right\rangle$ (23)
Here we have introduced core states from the atomic calculations $\left|\phi_{c,i}\right\rangle$. Notice that this is a different expression then the one with regular PAW in particular there are cross terms now between $\alpha=1$ and $\alpha>1$. In particular works for
$\hat{A}=\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right),\>\hat{A}=\frac{-\nabla^{2}}{2m}$ (24)
which are important for ground state energy determination. That is for the density and kinetic energy terms.
IV.2 Main modification to KS equations
We note that:
$\mathcal{T}_{M}^{\dagger}\hat{H}_{KS}\mathcal{T}_{M}\left|\tilde{\psi}_{n}\right\rangle=\varepsilon_{n}\mathcal{T}_{M}^{\dagger}\mathcal{T}_{M}\left|\tilde{\psi}_{n}\right\rangle$ (25)
Furthermore:
$S_{M}\equiv\mathcal{T}_{M}^{\dagger}\mathcal{T}_{M}=1+\sum_{\mathbf{R}}S_{\mathbf{R}}^{M}+\sum_{\mathbf{R}}S_{\mathbf{R}}^{M\dagger}+\sum_{\mathbf{R}}S_{\mathbf{R}}^{M\dagger}S_{\mathbf{R}}^{M}$
$=1+\left[\sum_{i\alpha\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle\left\langle\tilde{\phi}_{i\alpha}\right|\right]\sum_{i,\alpha\in\mathbf{R}}\left[\left|\phi_{j\beta}\right\rangle-\left|\tilde{\phi}_{j\beta}\right\rangle\right]\left\langle\tilde{p}_{j\beta}\right|+$
$+\sum_{i,\alpha\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle\left[\left\langle\phi_{i\alpha}\right|-\left\langle\tilde{\phi}_{i\alpha}\right|\right]\left[\sum_{j\beta\in\mathbf{R}}\left|\tilde{\phi}_{j\beta}\right\rangle\left\langle\tilde{p}_{j\beta}\right|\right]+$
$+\sum_{i,\alpha\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle\left[\left\langle\phi_{i\alpha}\right|-\left\langle\tilde{\phi}_{i\alpha}\right|\right]\sum_{j\beta\in\mathbf{R}}\left[\left|\phi_{j\beta}\right\rangle-\left|\tilde{\phi}_{j\beta}\right\rangle\right]\left\langle\tilde{p}_{j\beta}\right|$
$=1+\sum_{i,\alpha\in\mathbf{R}}\sum_{j\beta\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle\left[\left\langle\phi_{i\alpha}\mid\phi_{j\beta}\right\rangle-\left\langle\tilde{\phi}_{i\alpha}\mid\tilde{\phi}_{j\beta}\right\rangle\right]\left\langle\tilde{p}_{j\beta}\right|$
$\equiv 1+\sum_{i,\alpha\in\mathbf{R}}\sum_{j\beta\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle\Delta_{i\alpha j\beta}\left\langle\tilde{p}_{j\beta}\right|\neq\mathcal{T}^{\dagger}\mathcal{T}=S$ (26)
The reason for the inequality is that we have used a different (more accurate) resolution of identity to simplify the situation which leads to different results, in particular note again the cross terms between $\alpha=1$ and $\alpha>1$. Similarly for:
$\mathcal{T}_{M}^{\dagger}\hat{H}_{KS}\mathcal{T}_{M}=$
$=\sum_{i,\alpha\in\mathbf{R}}\sum_{j\beta\in\mathbf{R}}\left[\left\langle\phi_{i\alpha}\mid\hat{H}_{KS}\mid\phi_{j\beta}\right\rangle-\left\langle\tilde{\phi}_{i\alpha}\mid\hat{H}_{KS}\mid\tilde{\phi}_{j\beta}\right\rangle\right]\times$
×$\times\left|\tilde{p}_{i\alpha}\right\rangle\left\langle\tilde{p}_{j\beta}\right|+\hat{H}_{KS}$
$\equiv\hat{H}_{KS}+\sum_{i,\alpha\in\mathbf{R}}\sum_{j\beta\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle h_{i\alpha j\beta}\left\langle\tilde{p}_{j\beta}\right|\neq\mathcal{T}^{\dagger}\hat{H}_{KS}\mathcal{T}$ (27)
Again the resolution of identity changed the final result. Here $\hat{H}_{KS}=\nabla^{2}+V_{KS}$ where $V_{KS}$ is the KS potential. Furthermore the resolution of identity is used repeatedly for finding the ground state energy (see Section IV.3) and therefore also its derivative with respect to the electron density which is the value of the KS potential.
IV.3 Correlation and exchange and Hartree piece
We write for local correlation and exchange functionals such as Local Density Approximation (LDA) or Generalized Gradient Approximation (GGA):
$E_{XC}=E_{XC}\left(\tilde{n}\right)+\sum_{\mathbf{R}}E_{XC}\left(n_{\mathbf{R}}\right)-\sum_{\mathbf{R}}E_{XC}\left(\tilde{n}_{\mathbf{R}}\right)$ (28)
This is based on the idea that
$\tilde{n}=\tilde{n}_{\mathbf{R}}\>\left(\mathrm{in\>the\>sphere}\>\mathbf{R}\right)$ (29)
However for this we see that Eq. (19) is much better then Eq. (9) with explicit changes in the form of $\tilde{n}_{\mathbf{R}}$ seen from Eq. (23). Furthermore we now introduce the convenient notation:
$\left(f\mid g\right)=\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{f\left(\mathbf{r}_{1}\right)g\left(\mathbf{r}_{2}\right)}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}$ (30)
and
$\left(f\mid f\right)=\left(\left(f\right)\right)$ (31)
Similarly we have that the Coulomb piece is given by:
$E_{C}\left[n\right]$ $=U_{H}\left(\tilde{\rho}\right)$
$+\frac{1}{2}\sum_{\mathbf{\mathbf{R}}}\left[\left(\left(n_{\mathbf{R}}\right)\right)+2\left(n_{\mathbf{R}}\mid Z_{\mathbf{R}}\right)-\left(\left(\tilde{n}_{\mathbf{R}}+\tilde{Z}_{\mathbf{R}}\right)\right)\right]$ (32)
where
$U_{H}\left(\tilde{\rho}\right)=\frac{1}{2}\left(\left(\tilde{n}+\tilde{Z}_{\mathbf{R}}\right)\right)$ (33)
Where $Z_{\mathbf{R}}$ is the coulomb charge density and $\tilde{Z}_{\mathbf{R}}$ is the pseudized charge density satisfying the relationship:
$\int d^{3}\mathbf{r}\left|\mathbf{r}-\mathbf{R}\right|^{l}Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{R}}\right)\times$
×$\times\left(n_{\mathbf{R}}\left(\mathbf{r}\right)-\tilde{n}_{\mathbf{R}}\left(\mathbf{r}\right)+Z_{\mathbf{R}}\left(\mathbf{r}\right)-\tilde{Z}_{\mathbf{R}}\left(\mathbf{r}\right)\right)$
$=0$ (34)
Where we have repeatedly used Eq. (29) which is dependent on the resolution of unity (see Section IV.1) so Eq. (19) is much better then Eq. (9) for total energy calculations.
V Conclusion
In this work we have introduced a new version (extended basis set) of the PAW transformation (given in Eqs. (12), (13), (14) and (15)) which allows for an extended resolution of identity as part of the map between the pseudized problem and the all electron problem. This greatly improves the accuracy with which mapping reproduces the key PAW equation for expectation values local operators (given in Eq. (10)) as well as the various applications of that identity to exchange and correlation energies, Coulomb energies and the KS secular equation. This improves the accuracy of PAW at very limited computational cost and because of the mathematical form of our equations is nearly identical to that of regular PAW it can readily be incorporated into modern electronic structure implementations of PAW. In the future it would be of interest to include strong correlations into the calculations in the form of LDA+U or LDA+DMFT or LDA+GA [42, 43] with this new resolution of identity for the PAW method.